Calculate the period of a proton moving at 2.5 x 107 m/s perpendicularly to a 1.2T field of a cyclotron.

mass of a proton = 1.67 x 10-27 kg

charge on a proton = 1.602 x 10-19 C
Answer #1

The time period of a particle moving perpendicular to a field of a cyclotron, taking in relativistic considerations into account, is given by:

T=\frac{2\pi\gamma m}{Bq}\

where,it's given, m = 1.67 x 10-27 kg ; B = 1.2 T ; q = 1.602 x 10-19 C and (gamma) is Lorentz Factor.

Lorentz Factor is given by:

\gamma=\frac{1}{\sqrt{1-(\frac{v}{c})^2}}

where,it's given, v = 2.5 x 107 m/s and c = speed of light = 3 x 108 m/s [general fact]

The value of Lorentz Factor is :

\gamma=\frac{1}{\sqrt{1-(\frac{2.5\times10^7}{3\times10^8})^2}}=1.00349527178103

Therefore, the time period of the particle is:

gif.latex?T%3D%5Cfrac%7B2%5Ctimes3.14159

On simplification, we get:

T = 0.5248 ns

SUMMARY:

The time period of the proton particle is T=0.5248 ns